OpenStudy (anonymous):

For questions 5-7, use the following radical equation: radical x-1 = x-7 5. what would be the first step? a) Take the square root of each side to solve for x2 b) Square each side of the equation to remove the radicand c) Subtract 7 from each side d) Subtract 1 from each side 6. Factoring the equation will give the binomial: a) (x – 5)(x + 5) b)(x + 5)(x +10) c) (x – 5)(x – 10) d) None of the above 7. The correct solution for the radical equation is: a) 5 and 10 b) –5 and –10 c) 5 d)10

OpenStudy (anonymous):

a) square each side

OpenStudy (anonymous):

gives \[x-1=(x-7)^2=x^2-14x+49\]

OpenStudy (anonymous):

so \[x^2-15x+50=0\]